# Learning Zero-Shot Material States Segmentation, by Implanting Natural Image Patterns in Synthetic Data

### v5

@@ 6.1 Results on other benchmarks
The results of SAM, Materialistic, and MatSeg-trained nets on various benchmarks are given in table 2. These results support the conclusion of the previous section. The MatSeg-trained net outperformed other methods in cases of scattered segments, soft/gradual boundaries, and segmenting relatively similar states. Whether it is dust on roads, leaf diseases, corrosion, microscopy, or mud and dry ground. In contrast, SAM and Materialistic outperform MatSeg in benchmarks with materials that form bulk smooth boundaries, appear in single bulk shapes, have a strong correlation to objects, and don’t have gradual transitions (table 2). Materialistic is more similar to SAM in its results, despite the fact that both MatSeg and Materialistic are trained on synthetic data, while SAM is trained on real images. This supports the hypothesis that the data infusion method in MatSeg taps a fundamentally different distribution that was missed by both manual annotations and synthetic data generation approaches used to train SAM and Materialistic. However, it should also be noted that the accuracy of all methods on these benchmarks is limited, suggesting that a significant gap remains in this field.
@@ 6.2 Conclusion
Segmenting material states is fundamental to understanding nearly every aspect of the physical world and has numerous downstream applications. This work offers the first benchmark and dataset that deal with this task in a general zero-shot manner. This allows, for the first time, to train nets specifically for this task as well as to evaluate existing general methods. The results of this evaluation show that even the top foundation models struggle with this task. This work also offers a general, unsupervised approach for extracting patterns and textures from images and infusing them into synthetic scenes. This allows the creation of large-scale synthetic data that captures much of the complexity and diversity of the real world with no human effort. Nets trained on this data significantly outperformed the top foundation models on the MatSeg benchmark and many related zero-shot benchmarks (without training on the datasets). This suggests that the infusion method offers a way to close a major gap between real-world and synthetic data, and tap into patterns distribution missed by leading data generation approaches. The benchmark, dataset, and 300,000 extracted textures and PBR materials have been made available (appendix A).
@@ Appendix C Appendix: 3D scene building
Once a map is generated (section 3.1), the 3D scene creation follows the standard methods described in previous works[21]: 1) Load random 3D objects[11, 19]. 2) Load random PBR or BSDF material for each region of the map and use the generated UV map to set the materials on the object surface. 3) Load a panoramic HDRI image to act as a background and illumination source[3]. 4) Add random objects, a ground plane, and sometimes random light sources to create shadows and occlusion in the scene. 5) Set a random camera position and render the image. Scene generation was performed in Blender 4.0[8] with the Cycles rendering engine. The generation code and the dataset have been made available (appendix A).
@@ Appendix G Net and Training
To test the MatSeg dataset, we use it to train a net for class-agnostic soft material segmentation. Previous works have shown that representing materials in images as vector embeddings and using cosine similarity between these vectors to predict how similar the materials are is an effective method for matching materials states[21]. This representation also has the advantage that it gives a soft similarity, which is good for representing mixtures, transition states, and partial similarities. To apply this approach to image segmentation, we use the Unet-style neural net, which predicts 128 long descriptors for each pixel in the image (128-layers output map). This 128 vector represents the material for each pixel (fig. 6). The cosine similarity between the vectors of two pixels in the map gives the similarity of the materials in these two pixels. The average vector in a given region (group of pixels) belonging to the same material gives the average vector embedding for this specific material.
@@ G.1 Training
The training procedure is depicted in fig. 6 and goes as follows: a) A training batch contains a few randomly sampled images from the MatSeg dataset. b) The images are passed through the net to produce an embedding map with a vector per pixel. These vectors are normalized using L2 normalization. c) The predicted vectors of all pixels belonging to the same material are averaged to create an average vector per material (fig. 6). To determine which pixels belong to the same material, we use the Ground True (GT) mask of each material and take all the predicted vectors in this region from the descriptor map (fig. 6). e) Each of the average material vectors in the batch is then matched with each vector embedding in the predicted descriptor map. This gives a 2D cosine similarity map between each material in the batch and each pixel in the image. f) The similarity maps for a given image are stacked together and passed through a softmax (temperature of 0.2) to obtain the probability maps for the match between each pixel and each material. g) These probability maps, along with the GT one-hot masks, are used to calculate a cross-entropy loss. Note that the GT maps are turned from soft to hard (one-hot) segments by selecting for each pixel the GT material with the highest concentration (the highest value according to the GT map).

### v6

@@ 6.1 Results on other benchmarks
The results of SAM, Materialistic, and MatSeg-trained nets on various benchmarks are given in table˜2. These results support the conclusion of the previous section. The MatSeg-trained net outperformed other methods in cases of scattered segments, soft/gradual boundaries, and segmenting relatively similar states. Whether it is dust on roads, leaf diseases, corrosion, microscopy, or mud and dry ground. In contrast, SAM and Materialistic outperform MatSeg in benchmarks with materials that form bulk smooth boundaries, appear in single bulk shapes, have a strong correlation to objects, and don’t have gradual transitions (table˜2). Materialistic is more similar to SAM in its results, despite the fact that both MatSeg and Materialistic are trained on synthetic data, while SAM is trained on real images. This supports the hypothesis that the data infusion method in MatSeg taps a fundamentally different distribution that was missed by both manual annotations and synthetic data generation approaches used to train SAM and Materialistic. However, it should also be noted that the accuracy of all methods on these benchmarks is limited, suggesting that a significant gap remains in this field.
@@ 6.2 Conclusion
Segmenting material states is fundamental to understanding nearly every aspect of the physical world and has numerous downstream applications. This work offers the first benchmark and dataset that deal with this task in a general zero-shot manner. This allows, for the first time, to train nets specifically for this task as well as to evaluate existing general methods. The results of this evaluation show that even the top foundation models struggle with this task. This work also offers a general, unsupervised approach for extracting patterns and textures from images and infusing them into synthetic scenes. This allows the creation of large-scale synthetic data that captures much of the complexity and diversity of the real world with no human effort. Nets trained on this data significantly outperformed the top foundation models on the MatSeg benchmark and many related zero-shot benchmarks (without training on the datasets). This suggests that the infusion method offers a way to close a major gap between real-world and synthetic data, and tap into patterns distribution missed by leading data generation approaches. The benchmark, dataset, and 300,000 extracted textures and PBR materials have been made available (appendix˜A).
@@ Appendix C Appendix: 3D scene building
Once a map is generated (section˜3.1), the 3D scene creation follows the standard methods described in previous works[21]: 1) Load random 3D objects[11, 19]. 2) Load random PBR or BSDF material for each region of the map and use the generated UV map to set the materials on the object surface. 3) Load a panoramic HDRI image to act as a background and illumination source[3]. 4) Add random objects, a ground plane, and sometimes random light sources to create shadows and occlusion in the scene. 5) Set a random camera position and render the image. Scene generation was performed in Blender 4.0[8] with the Cycles rendering engine. The generation code and the dataset have been made available (appendix˜A).
@@ Appendix G Net and Training
To test the MatSeg dataset, we use it to train a net for class-agnostic soft material segmentation. Previous works have shown that representing materials in images as vector embeddings and using cosine similarity between these vectors to predict how similar the materials are is an effective method for matching materials states[21]. This representation also has the advantage that it gives a soft similarity, which is good for representing mixtures, transition states, and partial similarities. To apply this approach to image segmentation, we use the Unet-style neural net, which predicts 128 long descriptors for each pixel in the image (128-layers output map). This 128 vector represents the material for each pixel (fig.˜6). The cosine similarity between the vectors of two pixels in the map gives the similarity of the materials in these two pixels. The average vector in a given region (group of pixels) belonging to the same material gives the average vector embedding for this specific material.
@@ G.1 Training
The training procedure is depicted in fig.˜6 and goes as follows: a) A training batch contains a few randomly sampled images from the MatSeg dataset. b) The images are passed through the net to produce an embedding map with a vector per pixel. These vectors are normalized using L2 normalization. c) The predicted vectors of all pixels belonging to the same material are averaged to create an average vector per material (fig.˜6). To determine which pixels belong to the same material, we use the Ground True (GT) mask of each material and take all the predicted vectors in this region from the descriptor map (fig.˜6). e) Each of the average material vectors in the batch is then matched with each vector embedding in the predicted descriptor map. This gives a 2D cosine similarity map between each material in the batch and each pixel in the image. f) The similarity maps for a given image are stacked together and passed through a softmax (temperature of 0.2) to obtain the probability maps for the match between each pixel and each material. g) These probability maps, along with the GT one-hot masks, are used to calculate a cross-entropy loss. Note that the GT maps are turned from soft to hard (one-hot) segments by selecting for each pixel the GT material with the highest concentration (the highest value according to the GT map).